A bowl contains 6 blue balls, 8 red balls, 6 white balls, and 10 yellow balls. What is the probability of picking a yellow ball?


1
Expert's answer
2022-06-02T17:08:13-0400
"P(Y)=\dfrac{10}{6+8+6+10}=\dfrac{1}{3}"